# Axion Icebergs: Clockwork ALPs at hadron colliders

### v1

@@ Abstract
The conventional ultralight QCD axion is typically rendered invisible at collider experiments by its large decay constant. What could also hint at its possible existence is the observation of other (heavy) particles that are characteristically related to the light axion. One such scenario is afforded within the framework of the clockwork mechanism where the axion can have suppressed couplings with the gluons or photons while its companion axion-like particles (ALPs) have relatively unsuppressed couplings. We study a minimal clockwork model for the QCD axion invoking a KSVZ-like setup and examine the visibility of the ALPs $(a_{n})$ at the LHC through the process $pp\to a_{n}\,(+\,{\rm additional\,jets})$, $a_{n}\to\gamma\gamma$. The model contains $N$ ALPs with a decay constant $f$ and masses defined by a scale $m$ characteristic of the nearest-neighbour interactions of the scalar fields. For $10\lesssim m\lesssim 100$ GeV, $f\sim 1$ TeV and $N\sim\mathcal{O}(10)$, the full spectrum of ALPs is accessible and the corresponding diphoton invariant mass distribution comprises a unique signature of a wide band of resonances. For the case of light ALPs $(m\sim\mathcal{O}(10\,{\rm GeV}))$ with the axion being a dark matter candidate, the mass-splittings among the former are so small that the signal profile mimics that of a single broad resonance, or an axion iceberg. The effect subsides for heavier ALPs, albeit still exhibiting undulating peaks. For light ALPs, the scenario is imminently testable by the end of LHC’s Run 3 phase, with the estimated cumulative significance reaching the discovery threshold for an integrated luminosity of $\sim 300{\rm\,fb^{-1}}$. While the signals for the heavier ALPs in this minimal setup may not be as prominent within the ongoing LHC operation, one could expect to probe a wider parameter space of the model at the forthcoming HL-LHC.
1 Introduction
@@
The strong CP problem is, perhaps, one of the strongest motivations to look for new dynamics beyond the Standard Model (SM) of particle physics, in particular towards probing the existence of the light pseudoscalar particle predicted by the Peccei-Quinn (PQ) mechanism PQPhysRevLett ; WeinbergPhysRevLett ; WilczekPhysRevLett which is, arguably, the simplest and the most elegant solution to the problem proposed to date. The mechanism entails a pseudoscalar field $\phi$ (popularly called an axion or, more specifically, the QCD axion), that couples to gluons via the interaction term $(\phi/f)\tilde{G}G$. A shift in the field can, then, counter the CP violating topological term $(\theta\tilde{G}G)$ appearing in the SM as a result of the nontrivial structure of the QCD vacuum. There exist interesting alternatives to the PQ scenario as well, most notable among them are solutions based on a massless up quark Georgi:1981be ; Kaplan:1986ru ; Choi:1988sy ; Banks:1994yg ; Dine:2014dga and those invoking spontaneous CP violation (a la the Nelson-Barr mechanism) Nelson:1983zb ; Barr:1984qx . However, the $m_{u}=0$ solution within the standard theory currently stands in tension with the lattice QCD result of the topological mass contribution Alexandrou:2020bkd and models based on the Nelson-Barr mechanism are also subject to several theoretical challenges Dine:2015jga . Of course, the axion solution is no less prone to problems of its own. The quality problem and domain wall formation are some of the theoretical and cosmological issues that still plague the mechanism’s general viability, although there have been several noteworthy developments towards mitigating these issues Lazarides:1982tw ; Chatterjee:2019rch ; Ibe:2019yew ; Zhang:2023gfu . Nevertheless, the axion solution offers an interesting, albeit important, domain to explore phenomenologically, more so because the axion can potentially also play the role of an ultralight dark matter candidate Duffy:2009ig ; Chadha-Day:2021szb .
The axion is typically expected to have a very large decay constant ($f\gtrsim\mathcal{O}(10^{10}\,\mbox{GeV})$) depending on its mass so as to be consistent with the current experimental limits on its effective couplings with photons (as also electrons and nucleons) ParticleDataGroup:2024cfk ; AxionLimits . Furthermore, in the minimal scenarios the QCD axion mass scales inversely with the decay constant and, hence, a large $f$ also implies a very light mass for the axion. The light mass and extremely feeble couplings render the minimal QCD axion nearly invisible to most of the current and future experimental probes. For such scenarios the sensitivities of collider experiments are, understandably, quite poor. Indeed, most current collider searches concentrate, instead, on axion-like particles that either have no role in the strong CP issue Bauer:2018uxu ; dEnterria:2021ljz ; Florez:2021zoo or have other sources (new or extended confining sectors) which raise their masses substantially Rubakov:1997vp ; Berezhiani:2000gh ; Fukuda:2015ana ; Hook:2019qoh ; Agrawal:2017ksf ; Gherghetta:2020keg . This raises the question as to whether the discovery of a masive axion-like particle would necessarily negate the existence of the light QCD axion. We address this by exploring a theoretical scenario wherein the light and nearly invisible QCD axion is naturally accompanied by heavy pseudoscalars having couplings with the gluons and photons large enough to be visible at hadron colliders. One such scenario is realized in the context of a warped 5D (Randall-Sundrum) geometry where the axion field lives in the bulk and the resulting 4D theory contains the invisible axion as the zero mode along with a tower of massive Kaluza-Klein states Flacke:2006ad .
In this work, we focus on a relatively new alternative based on the clockwork mechanism which was originally introduced in the context of relaxions Kaplan:2015fuy ; Choi:2015fiu and later extended to general field theories Giudice:2016yja . The mechanism primarily entails the generation of hierarchical couplings or mass scales through the localization of the lightest particle in the theory space lattice defined by $N$ fields which interact with each other through nearest-neighbour couplings with a strength characterized by a parameter $q$. The massive modes, on the other hand, would be delocalized over the entire lattice. With a clockwork theory of $N+1$ pseudoscalars, then, the zero mode would correspond to the invisible axion with an exponentially large decay constant towards one end of the lattice compared to that of the $N$ delocalized heavy axions in the spectrum. The unsuppressed couplings of the heavy excitations raises the possibility that these may be produced at the LHC in numbers sufficient enough to be detectable. As it turns out, the individual heavy resonance continues to be barely visible over the background. However, their closeness in mass (a consequence of the clockwork paradigm) results in an intriguing signal at the LHC. A similar model has been discussed before in Farina:2016tgd , although with the primary focus on enhancing the photon coupling and studying the attendant phenomenology of the light axion, and in Higaki:2015jag in the context of the now obsolete $750$ GeV diphoton excess..
@@
The structure of the paper is as follows. In section, 2 we define a minimal clockwork model for a KSVZ-like QCD axion and briefly comment on how it addresses the strong CP problem. In addition to the light QCD axion, the model comprises a characteristic spectrum of massive axion-like particles (ALPs). We identify, in section 3, three benchmark points of the model suitable for probing the ALPs through the gluon-fusion channel at the LHC. In section 4, we discuss the corresponding signal and background profiles for $\sqrt{s}=13$ TeV and an integrated luminosity $\mathcal{L}=138\,{\rm fb}^{-1}$. Section 5, on the other hand, briefly describes the implications of the vector-like quark and the heavy radial scalars present in the model. Finally, we summarize the results and conclude in section 6.
2 The model
@@
We begin by constructing a clockwork realization of the minimal QCD axion, adopting a KSVZ-like Kim:1979if ; Shifman:1979if scenario. In the ensuing, we first detail the structure and spectrum of the clockwork sector and then proceed towards describing how the KSVZ setup can be implemented in the context of the clockwork mechanism.
2.1 Clockwork scalars
@@
$\begin{array}[]{rcl}{\cal L}_{\rm CW}&=&\displaystyle{\cal L}_{1}+{\cal L}_{2}\\[6.45831pt] {\cal L}_{1}&\equiv&\displaystyle\sum_{j=0}^{N}\left[(\partial_{\mu}\Phi_{j}^{\dagger})(\partial^{\mu}\Phi_{j})-\lambda\left(\Phi_{j}^{\dagger}\Phi_{j}-f^{2}\right)^{2}\right].\\[10.76385pt] {\cal L}_{2}&\equiv&\lambda^{\prime}\Lambda^{3-q}\sum_{j=0}^{N-1}\Phi_{j}^{\dagger}\Phi_{j+1}^{q}+\mbox{h.c.}\ ,\end{array}$ (1)
where $\lambda,\lambda^{\prime}$ are dimensionless real constants and $\Lambda\,(\ll f)$ is a characteristic scale associated with ${\cal L}_{2}$ (the exponent $q$ being, as yet, unspecified). In the absence of ${\cal L}_{2}$ (i.e., in the limit $\lambda^{\prime}\to 0$), the Lagrangian has a $U(1)^{N+1}$ global symmetry, the spontaneous breaking of which, at the scale $f$ (through $\langle\Phi_{j}^{\dagger}\Phi_{j}\rangle=f^{2}$ for all $j$) would lead to $(N+1)$ Goldstone bosons $\pi_{j}$, with the complex scalars being representable as
@@
The large global symmetry, discrete or continuous (and the attendant plethora of Goldstones in the broken phase) is neither well-motivated nor desirable. Clearly, it can be broken by arbitrary terms expressing interactions between the scalars. On the other hand, at least a single $U(1)$ needs to be present [^2] Understandably, an exact global symmetry is unrealistic because even in the absence of new symmetry breaking effects one would expect the global $U(1)$ to be explicitly broken by quantum gravity effects. The Planck suppressed symmetry violating terms thus generated can, in general, have serious implications for the QCD axion potential to be discussed later Kamionkowski:1992mf . There are, however, several interesting ways of increasing the degree of Planck suppression, e.g. by considering the scalars to be composites of entities belonging to a new confining sector Randall:1992ut . so that its spontaneous breaking could lead to the axion. The interaction Lagrangian ${\cal L}_{2}$ serves exactly this purpose. Representing perhaps the simplest set of operators which preserve a single $U(1)$ while exhibiting the clockwork mechanism, it can be understood in terms of nearest-neighbour interaction terms [^3] In general, one could also have non-nearest neighbour interactions while keeping the CW symmetry intact, an example being a term of the form $\Phi^{\dagger}_{j}\Phi^{q/p}_{j+1}...\Phi^{q^{p}/p}_{j+p}$. With reference to eq.(5) such a term adds, to the average mass scale of the pNGBs, a contribution $m_{a}^{(p)}\sim m(q^{p}/p)$ Ben-Dayan:2017rvr . For the case $q=2$ (which would be our primary choice as a benchmark for the collider analysis to be discussed) renormalizability necessitates $p=1,2$ for which the overall pNGB mass scale changes only slightly with $m_{a}\sim\sqrt{2}m_{a}^{(1)}$. The case for $q=3$ is a little different with $m_{a}\sim 3m_{a}^{(1)}$, although the qualitative aspects of our results and conclusions in this work would be applicable just as well. that exhibit locality in a theory space and, hence, the lattice defined by the fields $\Phi_{j}$ (each specifying a site in the theory space) can be regarded as the fifth dimension, although a continuum limit of such a construction is difficult to realize. As for the exponent $q$ in ${\cal L}_{2}$, a non-integer value would imply compositeness, thereby adding a further layer of complications. The clockwork mechanism, which will be employed in the following discussions, necessitates $q>1$. On the other hand, for $q>3$ one would obtain nonrenormalizable operators in $\mathcal{L}_{CW}$ implying that the explicit symmetry breaking takes place due to new dynamics operational at the heavy scale $\Lambda\gg f$ and beyond. This, however, is inconsistent with the notion that $U(1)^{N+1}$ is spontaneously broken at the lower scale $f$. To avoid such issues and for simplicity we consider a renormalizable theory with $1<q\leq 3$.
@@
with $\mathcal{Q}_{j}$ being the generators (charges) corresponding to the individual $U(1)_{j}$’s. Consequently, only one combination of the $\pi_{j}$’s would now be a true Goldstone boson, while all the rest would gain masses much smaller than $f$. In other words, the latter are now pseudo-Nambu-Goldstone bosons (pNGBs). To be specific, the full scalar Lagrangian, in the broken phase, can be written as,
$\begin{array}[]{rcl}\displaystyle\mathcal{L}^{Full}_{\Phi}&=&\displaystyle\sum_{j=0}^{N}\Big{[}\frac{1}{2}\partial_{\mu}\phi_{j}\partial^{\mu}\phi_{j}-\frac{1}{4}\lambda\left(\phi_{j}+f\right)^{4}\Big{]}\\[8.61108pt] &+&\displaystyle 2^{(1-q)/2}\lambda^{\prime}\Lambda^{3-q}\sum_{j=0}^{N-1}(\phi_{j}+f)(\phi_{j+1}+f)^{q}\cos{\frac{\pi_{j}-q\pi_{j+1}}{f}}\,.\end{array}$ (4)
@@
$C_{0j}=\mathcal{N}_{0}q^{-j}\,,\quad C_{nj}=\mathcal{N}_{n}\left[q\sin{\frac{jn\pi}{N+1}}-\sin\frac{(j+1)n\pi}{N+1}\right].$ (6)
Here $\mathcal{N}_{0}$ and $\mathcal{N}_{n}$ are the normalization factors,
@@
It is evident from eq.(6) that while the heavy pseudoscalars are nearly delocalized over the lattice, the massless mode — which will subsequently assume the role of the QCD axion $(a_{0})$ — is localized towards the $j=0$ site, i.e. it has an exponentially suppressed overlap with $\pi_{N}$ by virtue of the clockwork mechanism. Interestingly, this feature will eventually ensure that the light axion has a hierarchically large decay constant [^4] Note that the usage of the terminology “decay constant” here is different from that encountered in the context of processes such as $\pi^{\pm}\to l^{\pm}\nu$, where $\pi^{\pm}$ are the charged QCD pions. as compared to the heavy axions in the spectrum.
2.2 A KSVZ axion
Now that we have a massless pseudoscalar in the clockwork spectrum, we would like to promote it to the status of a realistic QCD axion. This is most easily carried out by embedding the clockwork sector in any one of the usual QCD axion models. Owing to the singlet nature of the CW scalars the KSVZ construction is the simplest one. To this end, we introduce one generation of Weyl fermions $\Psi_{L},\Psi_{R}$ with the following nontrivial charges under the SM gauge group and the $N$-th (global) Abelian group [^5] While it is possible to have the $\psi_{L,R}$ charged under all the additional $U(1)$s, this only adds a layer of complexity without any qualitative changes. of the CW sector ($G_{SM}\times U(1)_{N}$),

### v2

@@ Abstract
Scenarios with multiple pseudoscalars are interesting as they usually tend to provide a framework to naturally realize a light axion with a large decay constant which has rich applications in cosmology, especially in the context of inflation and light dark matter physics. On the other hand, from a particle physics perspective, this facilitates a solution to the strong CP problem with a low Peccei-Quinn symmetry breaking scale. One such realization is afforded within the framework of the clockwork mechanism where the axion can have suppressed couplings with the gluons or photons while its companion axion-like particles (ALPs) have relatively unsuppressed couplings, thereby facilitating detectability. We study a minimal clockwork model for the QCD axion invoking a KSVZ-like setup and examine the visibility of its unique multi-ALP $(a_{n})$ signature at the LHC, the most sensitive channel being $pp\to a_{n}\,(+\,{\rm additional\,jets})$ followed by $a_{n}\to\gamma\gamma$. In congruence with the astrophysical and cosmological bounds for the axion, a striking feature emerges for the case of light ALPs $(m\sim\mathcal{O}(10\,{\rm GeV}))$ wherein the mass-splittings among the former are so small that the signal profile mimics that of a single broad resonance, or an axion iceberg. The scenario is found to be imminently testable by the end of LHC’s Run 3 phase for an integrated luminosity of $\sim 300{\rm\,fb^{-1}}$. A larger average ALP mass, on the other hand, results in multiple closely-spaced peaks with a characteristic signal profile, and would be expected to be seen at the forthcoming HL-LHC. Possible additional signals are also listed.
preprint: APS/123-QED
I Introduction
@@
The strong CP problem is, perhaps, one of the strongest motivations to look for new dynamics beyond the Standard Model (SM) of particle physics. Of particular interest has been establishing the existence of the light pseudoscalar particle — popularly called an axion or, more specifically, the QCD axion — predicted by the Peccei-Quinn (PQ) mechanism Peccei and Quinn (1977); Weinberg (1978); Wilczek (1978) which is, arguably, the simplest and the most elegant solution [^1] There exist interesting alternatives to the PQ scenario, most notable among them being solutions based on a massless up quark Georgi and McArthur (1981); Kaplan and Manohar (1986); Choi et al. (1988); Banks et al. (1994); Dine et al. (2015) and those invoking spontaneous CP violation (a la the Nelson-Barr mechanism) Nelson (1984); Barr (1984). However, the $m_{u}=0$ solution within the standard theory currently stands in tension with the lattice QCD result of the topological mass contribution Alexandrou et al. (2020). Similarly, models based on the Nelson-Barr mechanism are also subject to several theoretical challenges Dine and Draper (2015). to the problem proposed to date.
Some theoretical and cosmological issues such as the quality problem and domain wall formation do question the mechanism’s general viability, although there have been several noteworthy developments towards mitigating these issues Lazarides and Shafi (1982); Chatterjee et al. (2020); Ibe et al. (2020); Zhang (2024). Nevertheless, the axion solution offers an interesting and important domain to explore phenomenologically, more so because the axion can potentially also play the role of an ultralight dark matter candidate Duffy and van Bibber (2009); Chadha-Day et al. (2022).
The axion is typically expected to have a very large decay constant ($f\gtrsim\mathcal{O}(10^{10}\,\mbox{GeV})$) —with the lower limit depending on its mass—so as to be consistent with the current limits (drawn from astrophysical and cosmological observations as well as from direct search experiments) on its effective couplings with photons (as also electrons and nucleons) Navas et al. (2024); O’Hare (2020). In other words, this essentially mandates that the PQ symmetry breaking scale ($f_{PQ}$) must be hierarchically larger than the electroweak (EW) scale. Furthermore, the QCD axion mass typically scales inversely with the decay constant and, hence, a large $f$ also implies a very light mass for the axion. The extremely large value of $f_{PQ}$ (and, the attendant smallness of the axion mass), thus, introduces a new hierarchy problem.
In addition, the tiny mass of the minimal QCD axion and its extremely feeble couplings render it nearly invisible to most of the current and future experimental probes, especially in the context of high energy experiments. Consequently, recent collider searches have tended to concentrate, instead, either on axion-like particles (ALPs), that have no role in the strong CP issue Bauer et al. (2019); d’Enterria (2021); Flórez et al. (2021), or on heavy QCD axions obtained by way of introducing new or extended confining sectors Rubakov (1997); Berezhiani et al. (2001); Fukuda et al. (2015); Hook et al. (2020); Agrawal and Howe (2018); Gherghetta et al. (2020) The latter case, although interesting in its own right, typically does not capture the cosmological attributes of a light axion, i.e. a role as a possible dark matter candidate and/or as the inflaton field in the so-called natural inflation scenario Freese et al. (1990); Adams et al. (1993). In view of this, a question naturally arises as to whether the discovery of a heavy axion would necessarily negate the existence of the light QCD axion. This aspect assumes particular significance as there are hints of several local excesses in searches for such scalar/pseudoscalar resonances Aad et al. (2023a); CMS (2023); Aad et al. (2022); Sirunyan et al. (2019); Aad et al. (2024a); Schmieden (2021) without these being robust enough to merit a discovery claim.
Scenarios with multiple aligned axions Kim et al. (2005); Choi et al. (2014) could, in principle, address these issues without assuming a large hierarchy in the individual decay constants and yet allowing for a quasi-flat direction in the field space, thereby leading to an effective decay constant $f_{a}\gg f_{PQ}$ for the QCD axion. Ref.Higaki et al. (2016a) sought to exploit this with $f_{PQ}$ close to the EW scale and ALPs with masses $\gtrsim m_{EW}$.
In this work, we seek to present a theoretically better-motivated model that also leads to a very distinctive multi-ALP spectrum. With ALP masses $\lesssim m_{EW}$, this exposition would not only be complementary to the scenarios of refs.Higaki et al. (2016a); Farina et al. (2017) but would also confront features in existant data at the LHC Aad et al. (2023a); CMS (2023); Aad et al. (2022); Sirunyan et al. (2019); Aad et al. (2024a); Schmieden (2021). Furthermore, the light QCD axion inherent to the scenario, apart from being a viable DM candidate, would likely be testable at the current and/or next generation of dedicated experiments. To this end, we focus on the clockwork paradigm, originally proposed in refs.Choi et al. (2014); Choi and Im (2016); Kaplan and Rattazzi (2016); Giudice and McCullough (2017), as it automatically generates hierarchical couplings or mass scales through the localization of the lightest particle in the theory space lattice defined by $N$ fields which interact with each other through nearest-neighbour couplings with a strength characterized by an ${\cal O}(1)$ parameter $q$. The massive modes, on the other hand, are delocalized over the entire lattice.
To realize a QCD axion within such a construction, we consider a model with a clockwork sector comprising $N+1$ complex scalars $\Phi_{j}(j=0,...,N)$, each charged under a unique $U(1)$ symmetry. Inspired by the KSVZ model Kim (1979); Shifman et al. (1980), we introduce a pair of new, $SU(2)_{L}$ singlet, coloured Weyl fermions ($\Psi_{L,R}$), chiral under the $N$th group $U(1)_{N}$, which couple to the $N$-th complex scalar. With only a single species of new coloured fermions, the putative domain wall problem, on account of a spontaneous breaking of the discrete symmetry present in the QCD-induced potential for the axion, is well under control, and such a KSVZ-like model is free of the problems that often beset, say, a DFSZ-like Zhitnitsky (1980); Dine et al. (1981) scenario.
A spontaneous breaking of the full symmetry ($U(1)^{N+1}$) at a scale $f$ (an analogue of $f_{PQ}$) in the clockwork sector, alongwith an explicit breaking of the symmetry to a single factor $U(1)_{CW}$ due to the nearest-neighbour interactions characterized by an effective mass scale $m\ll f$, leads to one massless Nambu-Goldstone Boson (NGB) and $N$ pseudo-NGBs in the spectrum. The chiral anomaly, then, induces a coupling of the light NGB to the gluonic topological operator ($\tilde{G}G$) with an exponentially large effective decay constant $\sim q^{N}f$, thanks to the localization mechanism. Below the confinement scale, this leads to a small mass for the light NGB through nonperturbative effects. Therefore, for $q>1$ and a sizable number of CW fields, the light NGB can be identified with the QCD axion even for the fundamental SSB scale $f$ being as low as a few TeVs. The pNGBs, on the other hand, act as ALPs with masses of the order $\sim mq$ (thereby creating a mass gap) and couple to gluons with an effective decay constant comparable to the scale $f$. The nonzero hypercharge of $\Psi_{L,R}$ not only allows them to decay, but also engenders loop-induced couplings of the ALPs to a pair of photons and $Z$’s. At a hadron collider (current or future), ALPs are dominantly produced through gluon-gluon fusion, and their most prominent, yet unique, signature would be the occurrence of multiple closely spaced resonance peaks in the $\gamma\gamma$, $ZZ$ and $Z\gamma$ channels. The closeness in the ALP masses emerges naturally in the model as a consequence of the clockwork construction wherein the mass-splitting varies as $\sim 2m/N$, (with $N\gtrsim 10$ for adequate enhancement of the axion decay constant). Understandably, for lighter ALPs ($m\lesssim 100$ GeV), the mass-splittings would be small, and in certain cases so much so that the currently operating detectors at the LHC might not be able to completely resolve the individual peaks, leading to the appearance of the entire ALP spectrum as a single broad resonance. We argue that both signals (multiple closely spaced narrow peaks or a single anomalously broad peak) are unique to the clockwork paradigm and can be generated neither by a single, phenomenologically viable ALP candidate, nor by a generic model of multiple ALPs without invoking an ad hoc fine-tuning in the parameter space. Additionally, such signatures are complementary to the possibility of observing periodic (in invariant mass) signals in the case of relatively heavier ALPs, e.g. in the spirit of the analyses carried out in the context of the gravity model in refs.Giudice et al. (2018); Beauchesne and Kats (2020); Aad et al. (2023b). Driven by this notion, we assume benchmark points with ALP masses at or below the EW scale and investigate their detection prospects in the diphoton channel at the LHC. The most interesting outcome is found in the case of the mass parameter $m$ being small, where although the individual resonances are barely visible over the background, the significance of the conjoint signal profile, composed of multiple overlapping peaks, is projected to reach the discovery threshold even as early as by the end of the LHC’s Run 3 phase. An analysis of the diphoton final state, which is arguably the cleanest channel to probe the ALPs, is also motivated by the number of dedicated collider searches for light scalar resonances, near and below the EW scale, performed by the ATLAS Aad et al. (2023a, c) and CMS CMS (2023) collaborations in the recent years. Thus, our study also appeals to the case for further experimental probes of light scalar or pseudoscalar particles in the diphoton channel, scanning not only for one but multiple closely-spaced candidates or even in the form of a broad resonance.
Note that although the introduction of a clockwork sector does not alter or improve upon the basic features of the PQ mechanism, it does provide ways to mitigate certain issues present in the minimal QCD axion models, which in turn makes this avenue even more desirable to explore. For one, the possibility of a low scale $f_{PQ}$ helps improve the quality of the axion solution against putative quantum gravity effects [^2] Since the degree of the global symmetry breaking depends on the exact manifestation of quantum gravity effects near the Planck scale, a robust estimate of the quality problem in our model, given the complexity of the global symmetry involved, is difficult to obtain. Nevertheless, under the assumption that the leading QG effect arises as a dim-5 operator of the form $(g/M_{Pl})\left(|\Phi_{j}|^{2}\right)^{2}\Phi_{j^{\prime}}$ in the IR, the required fine-tuning for an acceptable PQ solution can be characterized in terms of an upper limit on the coupling $|g|$, which, in a model with $f_{PQ}\sim\mathcal{O}(1\,\mbox{TeV})$, can be estimated to be $|g|\lesssim 10^{-12}/N^{4}$. E.g., $N\sim 30$ implies $|g|\lesssim 10^{-18}$. Although this seems a very stringent upper limit in the absolute sense, it is nevertheless many orders of magnitude larger than what is required in the minimal models with $f_{PQ}\sim 10^{12}$ GeV, namely $|g|\lesssim 10^{-55}$ Kamionkowski and March-Russell (1992). Thus, even with the aforementioned ambiguities, it can be qualitatively argued that the clockwork scenario offers some respite from the severe quality issues of the typical axion models, if not a solution. The severity may be relaxed further with the imposition of additional discrete symmetries in the UV in order to restrict lower dimensional operators. However, we desist from doing so. Kamionkowski and March-Russell (1992); Higaki et al. (2016a). Secondly, the geometric progression of the light axion’s distribution over the lattice can be exploited to obtain an enhanced coupling with photons, independent of the coupling to the gluons, thereby opening a window in the parameter space that falls outside of the usual QCD band on the photon-coupling vs axion-mass plane. This aspect was studied in detail in ref.Farina et al. (2017).
@@
The structure of the paper is as follows. In section II, we define a minimal clockwork model for a KSVZ-like QCD axion and demonstrate how it addresses the strong CP problem. We identify, in section III, four benchmark points of the model suitable for probing the ALPs through the gluon-fusion channel at the LHC. In section IV, we discuss the corresponding signal and background profiles for $\sqrt{s}=13$ TeV and an integrated luminosity $\mathcal{L}=138\,{\rm fb}^{-1}$. Section V, on the other hand, briefly describes the implications of the vector-like quark and the heavy radial scalars present in the model. Finally, we summarize the results and conclude in section VI.
II The model
@@
We begin by constructing a clockwork realization of the minimal QCD axion, adopting a KSVZ-like Kim (1979); Shifman et al. (1980) scenario. In the ensuing, we first detail the structure and spectrum of the clockwork sector and then proceed towards describing how the KSVZ setup can be implemented in the context of the clockwork mechanism.
II.1 Clockwork scalars
@@
$\begin{array}[]{rcl}{\cal L}_{\rm CW}&=&\displaystyle{\cal L}_{1}+{\cal L}_{2}\,,\\[6.45831pt] {\cal L}_{1}&\equiv&\displaystyle\sum_{j=0}^{N}\left[(\partial_{\mu}\Phi_{j}^{\dagger})(\partial^{\mu}\Phi_{j})-\lambda\left(\Phi_{j}^{\dagger}\Phi_{j}-\frac{f^{2}}{2}\right)^{2}\right]\,,\\[10.76385pt] {\cal L}_{2}&\equiv&\lambda^{\prime}\Lambda^{3-q}\sum_{j=0}^{N-1}\Phi_{j}^{\dagger}\Phi_{j+1}^{q}+\mbox{h.c.}\ ,\end{array}$ (1)
where $\lambda,\lambda^{\prime}$ are dimensionless real constants and $\Lambda\,(\ll f)$ is a characteristic scale associated with ${\cal L}_{2}$ (the exponent $q$ being, as yet, unspecified). To understand better the pieces ${\cal L}_{1,2}$, let us examine them in isolation. In the absence of ${\cal L}_{2}$ (i.e., in the limit $\lambda^{\prime}\to 0$), the Lagrangian has a $U(1)^{N+1}$ global symmetry, the spontaneous breaking of which, at the scale $f$ (through $\langle\Phi_{j}^{\dagger}\Phi_{j}\rangle=f^{2}/2$ for all $j$) would lead to $(N+1)$ Goldstone bosons $\pi_{j}$, with the complex scalars being representable as
@@
The large global symmetry, discrete or continuous (and the attendant plethora of Goldstones in the broken phase) is neither well-motivated nor desirable. Clearly, it can be broken by arbitrary terms expressing interactions between the scalars. On the other hand, at least a single $U(1)$ needs to be present [^3] As previously mentioned, an exact global symmetry is unrealistic because even in the absence of new symmetry breaking effects one would expect the global $U(1)$ to be explicitly broken by quantum gravity effects. The Planck suppressed symmetry violating terms thus generated can, in general, upset the PQ solution through corrections to the QCD axion potential to be discussed later Kamionkowski and March-Russell (1992). While the small $f_{PQ}$ in our scenario already alleviates this so-called quality problem, it may be further assuaged by ensuring that the Planck suppressed operators in the theory appear only at a sufficiently high mass-dimension, e.g., by considering the CW scalars to be composites of entities belonging to new confining sector(s) Randall (1992); Coy et al. (2017). so that its spontaneous breaking could lead to the axion. The interaction Lagrangian ${\cal L}_{2}$ serves exactly this purpose [^4] The requirement $\Lambda\ll f$ at best introduces a little hierarchy and for $q=2$, as we would adopt, it represents only a soft breaking of the symmetry.. Representing perhaps the simplest set of operators which preserve a single $U(1)$ while exhibiting the clockwork mechanism, it can be understood in terms of nearest-neighbour interaction terms [^5] In general, one could also have non-nearest neighbour interactions while keeping the CW symmetry intact, an example being a term of the form $\Phi^{\dagger}_{j}\Phi^{q/p}_{j+1}...\Phi^{q^{p}/p}_{j+p}$. With reference to eq.(5) such a term adds, to the average mass scale of the pNGBs, a contribution $m_{a}^{(p)}\sim m(q^{p}/p)$ Ben-Dayan (2019). For the case $q=2$ (which would be our primary choice as a benchmark for the collider analysis to be discussed) renormalizability necessitates $p=1,2$ for which the overall pNGB mass scale changes only slightly with $m_{a}\sim\sqrt{2}m_{a}^{(1)}$. The case for $q=3$ is a little different with $m_{a}\sim 3m_{a}^{(1)}$, although the qualitative aspects of our results and conclusions in this work would be applicable just as well. that exhibit locality in a theory space and, hence, the lattice defined by the fields $\Phi_{j}$ (each specifying a site in the theory space) can be regarded as the fifth dimension [^6] A continuum theory that reproduces this theory on discretization, though, requires the introduction of certain additional features and we desist from discussing it. As for the exponent $q$ in ${\cal L}_{2}$, a non-integer value would imply compositeness, thereby adding a further layer of complications. The clockwork mechanism, which will be employed in the following discussions, necessitates $q>1$. On the other hand, for $q>3$ one would obtain nonrenormalizable operators in $\mathcal{L}_{CW}$ implying that the explicit symmetry breaking takes place due to new dynamics operational at the heavy scale $\Lambda\gg f$ and beyond. Furthermore, $q>3$ would also tend to destabilize the vacuum. To avoid such issues and for simplicity we consider a renormalizable theory with $1<q\leq 3$.
@@
with $\mathcal{Q}_{j}$ being the generators (charges) corresponding to the individual $U(1)_{j}$’s. Consequently, only one combination of the $\pi_{j}$’s would now be a true Goldstone boson, while all the rest would gain masses much smaller than $f$. In other words, the latter are now only pNGBs. To be specific, the full spin-0 Lagrangian, in the broken phase, can be written as,
$\begin{array}[]{rcl}\mathcal{L}^{Full}_{\Phi}=\sum_{j=0}^{N}\Big{[}\frac{1}{2}\partial_{\mu}\phi_{j}\partial^{\mu}\phi_{j}-\frac{1}{4}\lambda\left(\phi_{j}+f\right)^{4}\Big{]}+2^{(1-q)/2}\lambda^{\prime}\Lambda^{3-q}\displaystyle\sum_{j=0}^{N-1}(\phi_{j}+f)(\phi_{j+1}+f)^{q}\cos{\frac{\pi_{j}-q\pi_{j+1}}{f}}\,.\end{array}$ (4)
@@
$\begin{split}C_{0j}&=\mathcal{N}_{0}q^{-j}\,,\\ C_{nj}&=\mathcal{N}_{n}\left[q\sin{\frac{jn\pi}{N+1}}-\sin\frac{(j+1)n\pi}{N+1}\right].\end{split}$ (6)
Here $\mathcal{N}_{0}$ and $\mathcal{N}_{n}$ are the normalization factors
@@
It is evident from eq.(6) that while the heavy pseudoscalars are nearly delocalized over the entire lattice, the massless mode — which will subsequently assume the role of the QCD axion $(a_{0})$ — is localized towards the $j=0$ site, i.e. it has an exponentially suppressed overlap with $\pi_{N}$ by virtue of the clockwork mechanism. Interestingly, this feature will eventually ensure that the light axion has a hierarchically large decay constant [^7] Note that the usage of the terminology “decay constant” here is different from that encountered in the context of processes such as $\pi^{\pm}\to l^{\pm}\nu$, where $\pi^{\pm}$ are the charged QCD pions. as compared to the heavy axions in the spectrum.
II.2 A KSVZ axion
Now that we have a massless pseudoscalar in the clockwork spectrum, we would like to promote it to the status of a realistic QCD axion. This is most easily carried out by embedding the clockwork sector in any one of the usual QCD axion models. Owing to the singlet nature of the CW scalars, the KSVZ construction is the simplest one. To this end, we introduce one generation of Weyl fermions $\Psi_{L},\Psi_{R}$ with the following nontrivial charges under the SM gauge group and the $N$-th (global) Abelian group [^8] The chiral configuration only at the $N$-site ensures that the color anomaly is localized at a single site which enables us to exploit the CW mechanism to generate a large effective decay constant for the axion, as we will see shortly. On the other hand, while it is possible to have the $\Psi_{L,R}$ charged nonchirally under all the additional $U(1)$s, this only adds a layer of complexity without any qualitative changes in the phenomenology. of the CW sector ($G_{SM}\times U(1)_{N}$),